OpenStudy (anonymous):

9. What is the measure of one of the interior angles in a regular convex hexagon?

OpenStudy (anonymous):

120 degrees

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

ok - oh i remembered the formula (180(n-2))/n so for a 6 sided figure - a hexagon, it = (180(6-2))/6 = 720 / 6 = 120 degrees this is one of the interior angle s in a regular hexagon